OpenStudy (anonymous):

What are the foci of the ellipse? Graph the ellipse. 18x2 + 36y2 = 648

OpenStudy (anonymous):

@genius12

OpenStudy (anonymous):

@jim_thompson5910 @Directrix @.Sam. @Luis_Rivera can anybody help me please

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

If you divide both sides by 648, you get the following: 1/36x^2 + 1/18y^2 = 1 The foci are at 3sqrt(2) and -3sqrt(2)

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

@genius12 Is it +- first or 0

OpenStudy (anonymous):

This is what i mean because i know its between those to.

OpenStudy (anonymous):

@genius12 Help please

OpenStudy (sirm3d):

because x^2 has the bigger denominator, the foci lie along the x-axis

OpenStudy (anonymous):

What do you mean? The foci are achieved through his equation: \[a^2-c^2=b^2\] Where a^2 in this is case is 36, c^2 we need to solve for, and b^2 is 18. This gives: \[36-c^2=18\rightarrow c = \pm \sqrt18=\pm 3\sqrt2\]

OpenStudy (anonymous):

I got my foci already. It's between 2 answers that i posted above. Do you know which one ?

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

Thank you @genius12 :)

OpenStudy (sirm3d):